A ladder 10 feet long leans against a vertical wall, touching it 8 feet above the ground. What
is the slope of the ladder? Can a person 6-feet tall pass under the ladder 1 feet away from
the wall? Can the same person pass under the ladder 2 feet away from the wall? Use the
concept of slopes for your solution.

Answer by Alan3354(69443)

A ladder 10 feet long leans against a vertical wall, touching it 8 feet above the ground. What is the slope of the ladder?
Can a person 6-feet tall pass under the ladder 1 feet away from the wall?
Can the same person pass under the ladder 2 feet away from the wall?
----------------
It's a 6-8-10 right triangle.
With the intersection of the ground and the wall at the Origin, the ladder has a slope of -8/6 = -4/3
---
The equation of "the ladder" is y-8 = (-4/3)x
y = -4x/3 + 8
==============================
At x=1 (1 foot from the wall) y = 7 2/3 feet so a 6-foot person would clear.
At x=2, y = 5' 4"